OpenStudy (anonymous):

trying to verify 1/(1*3*5*...*(2n-5)) = ((2^n)(2n-3)(2n-1)(n!))/(2n)! never seen a similar factorial problem

OpenStudy (anonymous):

So are you supposed to prove it's an identity? Or establish that it is an identity?

OpenStudy (anonymous):

I think the right side is supposed to reduce to look like the left side

OpenStudy (anonymous):

What I mean is, do you have to show that the identity is true, or do you have to rewrite/simplify one side so that it matches the other exactly? The first would involve a proof by induction, and the other could go any one of a number of ways.

OpenStudy (anonymous):

unfortunately the only instructions I have on this problem is "Verify the formula: "

OpenStudy (anonymous):

This should get you started:\[\begin{align*} \frac{1}{1\times3\times5\times\cdots\times(2n-7)\times(2n-5)}&=\frac{2\times4\times\cdots\times(2n-6)\times(2n-4)}{(2n-5)!}\\ &=\frac{2(1)\times2(2)\times\cdots\times2(n-3)\times2(n-2)}{(2n-5)!}\\ &=\frac{2^{n-2}\bigg[1\times2\times\cdots\times(n-3)\times(n-2)\bigg]}{(2n-5)!}\\ &=\frac{2^{n-2}(n-2)!}{(2n-5)!} \end{align*}\] That's what you get just from working with the left side. To get what you're given on the right side, you would just multiply by whatever you need. (made a small mistake)

OpenStudy (anonymous):

Thanks! :-) I'm usually good at simplifying factorials, but this one has been a beast

OpenStudy (anonymous):

yw